the angle measures of a triangle are 2x+3 degrees, 9x-7 degrees, and 7x+4 degrees.which of the following is NOT one of the angle measures of the triangle A)23 degrees B)74 degrees C)83 degrees D)97 degrees
1
Expert's answer
2020-09-17T13:24:09-0400

We have three angles:


"\alpha=2x+3,\\
\beta=9x-7,\\
\gamma=7x+4."


The sum of angles in a flat 2D triangle is 180°:


"180°=\alpha+\beta+\gamma,\\
180°=(2x+3)+(9x-7)+(7x+4),\\
180°=18x,\\
x=10°."

Substituting this value for any of the given angles, we see that 97° is what does not belong to this triangle.